OpenStudy (anonymous):

a pure hydrocarbon gas is burned (combustion) to yield 2.641g of carbon dioxide and 1.442g of water. what is the empirical and molecular formulas.

OpenStudy (welshfella):

first work out the amount of carbon in 2.641 g of carbon dioxide and the amount of hydrogen in 1.442 g of water

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok do i need to find their molar masses?

OpenStudy (welshfella):

use the relative atomic masses C = 12, O = 16 and H = 1.008

OpenStudy (anonymous):

Well yeah, but I am asking f I need to find their molar masses first?

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

Well I know that in CO2 there are 32 g of carbon dioxide, and in H2O 2.016

OpenStudy (anonymous):

Right?

OpenStudy (welshfella):

fraction of carbon on CO2 = 12 / (12+32)

OpenStudy (anonymous):

Wait why would I divide the 12 by the 12+32?

OpenStudy (welshfella):

and H in H2O is (2.016) / 18.016

OpenStudy (welshfella):

1 C + 2 oxygens = 12 + 2*16

OpenStudy (anonymous):

What formula would I need to use? Or what formula does this apply to

OpenStudy (anonymous):

Oh ok, that would be the molar mass for co2

OpenStudy (anonymous):

Right?

OpenStudy (welshfella):

carbon dioxxide = CO2 and water = H2O

OpenStudy (welshfella):

Oh OK thats what you call the molar mass

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

So I find the molar masses, and then what?

OpenStudy (welshfella):

so multiplying those fractions by 2.641 and 1.442 will give the amounts of C and H

OpenStudy (anonymous):

Ok, and from that how could I calculate my empirical formulas? Do I need to convert my answers to %s for % comp?

OpenStudy (welshfella):

No you get the ratio of C to H in the hydrocarbon by dividing these 2 masses by 12 for C and 1.008 for H.

OpenStudy (anonymous):

OH!!!

OpenStudy (welshfella):

i get the 2 masses as 0.72027 for C and 0,16136 for H.

OpenStudy (anonymous):

Wait, how did you get those numbers? Sorry I got a bit lost

OpenStudy (welshfella):

divide the first by 12 and the second by 1.008 and you'l get the ratio C:H

OpenStudy (welshfella):

C = ( 12 / 44 )* 2.641

OpenStudy (anonymous):

oh so 12/44?

OpenStudy (welshfella):

yea

OpenStudy (welshfella):

work that out and it should come to 0.72027 g

OpenStudy (anonymous):

Ok, wait. Could you pleas go over the whole thing again? I don't know why but I can;t see anything past the last post

OpenStudy (anonymous):

Sorry, and thank you

OpenStudy (welshfella):

Burning the hydrocarbon produces CO2 and water (H2O). You need to find the masses of C and H in these 2 compounds. fraction of C in CO2 = 12 (12 + 32) ( thats the molar mass of carbon / molar mass of CO2) similar thing for H in H2O

OpenStudy (anonymous):

Ok, could I solve really quickly and then get back to you with an answer?

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

Ok im working on it

OpenStudy (anonymous):

Ok wait is the amount of c in co2 .2727?

OpenStudy (anonymous):

And would the amount of h in h2o be .119? @welshfella

OpenStudy (welshfella):

C is 0.2727 H is 0.1119

OpenStudy (anonymous):

yeahI got that with the 12/12+32 equations

OpenStudy (anonymous):

Is that right?

OpenStudy (welshfella):

yea now you need to use these to find the amounts of C and H in the products of the reaction

OpenStudy (anonymous):

How would I do that? Do I divide by the 2.641g and the 1.442g?

OpenStudy (welshfella):

multiply

OpenStudy (anonymous):

Oh.. Really I would multiply?

OpenStudy (welshfella):

0.2727 * 2.641 0.119 * 1.442

OpenStudy (welshfella):

* correction 0.1119 * 1.442

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So for the first one the carbon I got .72

OpenStudy (anonymous):

Would that be correct?

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

Ok and for the hydrogen I got .17

OpenStudy (welshfella):

0.16

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Now what would I do with these two numbers?

OpenStudy (welshfella):

now divide by the molar masses 0.72 by 12 and 0.16 by 1.008

OpenStudy (anonymous):

Oh, ok

OpenStudy (welshfella):

this will give you the ratio of C to H in the hydrocarbon

OpenStudy (anonymous):

I got .06 and .15

OpenStudy (anonymous):

Is this right?

OpenStudy (welshfella):

0.16 is closer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah it would be .158 so we just round it up right?

OpenStudy (welshfella):

so ratio of C:H is 6:16 somolecular formula is C6H16

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and then I think i now how to find the molecular formula.. I'll get back to you with that in a minute

OpenStudy (welshfella):

Do you have the molecular mass of the hydrocarbon?

OpenStudy (anonymous):

Not yet

OpenStudy (anonymous):

What formula would I need to use? n=PV/RT

OpenStudy (anonymous):

Wait no never mind

OpenStudy (welshfella):

well to find the molecular formular we would need the vapour density of the gas Mocelar mass = 2 * vapour density all we have now is the formula C6H16 which might be its molecular mass the empirical foemula could be propane C3H8 C6H16 could be a conjugate of propane

OpenStudy (anonymous):

I need the molar mass and the empirical formula to find the molecular formula right?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

but we havent got molar mass

OpenStudy (welshfella):

if it had been given in the question we could be sure

OpenStudy (welshfella):

i would go with C3H8 for the empirical formula and C6H16 for the molecular formular

OpenStudy (welshfella):

gotta go right now

OpenStudy (anonymous):

Ok! Thank you!